tyson:

Marnie conducted a survey of the cats in four shelters and wrote the following observations: 4 of the 18 cats in Shelter 1 were black 13 of the 24 cats in Shelter 2 were black 12 of the 20 cats in Shelter 3 were black 18 of the 25 cats in Shelter 4 were black Based on the survey, which shelter had the greatest percentage of black cats? (5 points) Shelter 1 Shelter 2 Shelter 3 Shelter 4

Elsa213:

What do you think?

tyson:

shelter 2?

AngeI:

So for greatest percentage you could normally take the number of black cats and divide by the total number of cats like 4/18 But this one you can more or less eyeball it lol

tyson:

its shelter 4

Elsa213:

4/18=22% 13/24=54% 12/20=60% 18/25=72%

tyson:

right?

Elsa213:

I agree.

adovbush:

Yup

tyson:

ok thx

adovbush:

I wish I had these kind of questions on my quizzes :/